# Multi-Bubble Blow-up Analysis for an Almost Critical Problem

### v1

@@
Abstract: Consider a smooth, bounded domain $\Omega\subset\mathbb{R}^{n}$ with $n\geq 4$ and a smooth positive function $V$. We analyze the asymptotic behavior of a sequence of positive solutions $u_{\varepsilon}$ to the equation $-\Delta u+V(x)u=u^{\frac{n+2}{n-2}-\varepsilon}$ in $\Omega$ with zero Dirichlet boundary conditions, as $\varepsilon\to 0$. We determine the precise blow-up rate and characterize the locations of interior concentration points in the general case of multiple blow-up, providing an exhaustive description of interior blow-up phenomena of this equation. Our result is established through a delicate analysis of the gradient of the corresponding Euler-Lagrange functional.
@@ 1 Introduction and Main Results
where $\Omega$ is a smooth bounded domain in $\mathbb{R}^{n}$, $n\geq 4$, the potential $V$ is a $C^{3}$ positive function on $\overline{\Omega}$, $p+1=(2n)/(n-2)$ is the critical Sobolev exponent for the embedding $H^{1}_{0}(\Omega)\hookrightarrow L^{q}(\Omega)$ and $\varepsilon$ is a small positive parameter.
@@
Over the past decades, a substantial body of literature has developed on the blow-up phenomena of solutions to semilinear equations with critical exponents, a topic too broad to summarize fully. Here, we highlight only a selection of some works, from which we hope a more comprehensive bibliography can be constructed. In the case where $V\equiv 0$, the asymptotic behavior of solutions that blow up at a single point has been well-established in [12, 33, 24]. The scenario of multiple concentration points, still for $V\equiv 0$, was later investigated in [34, 7, 35]. For $V\not\equiv 0$, single-blow-up behavior in the interior has been studied in the three-dimensional case by [23]. The focus of this paper is on multiple blow-up points in the interior for dimensions greater than or equal to 4. Other blow-up scenarios have also been studied, including those for sign-changing solutions and for positive solutions under Neumann boundary conditions. For additional references, we refer the reader to [26, 19, 22, 29, 1, 2, 27, 14, 25] and the works cited therein. For the inverse problem of constructing blow-up solutions in this setting, we direct the reader to [3, 10, 8, 7, 20, 11, 36, 37, 17, 30, 32] and their respective references.
@@
[…]
@@
Since $V$ is a continuous positive function on $\overline{\Omega}$, we see that this norm is equivalent to the two norms ||.$||.||_{0}$ and ||.$||.||_{1}$ of $H^{1}_{0}(\Omega)$ and $H^{1}(\Omega)$ respectively.
Note that all solutions $u_{\varepsilon}$ of $(\mathcal{P}_{\varepsilon,V})$ satisfy $\|u_{\varepsilon}\|\geq C$ with a positive constant $C$ independent of the parameter $\varepsilon$. Therefore, the concentration compactness principle ([38, 28]) implies that if $u_{\varepsilon}$ is an energy-bounded solution of $(\mathcal{P}_{\varepsilon,V})$ that weakly converges to 00, then $u_{\varepsilon}$ must concentrate at a finite number $N$ of points in $\overline{\Omega}$. More precisely, $u_{\varepsilon}$ can be expressed as:
$u_{\varepsilon}=\sum_{i=1}^{N}\delta_{{a_{i,\varepsilon}},\lambda_{i,\varepsilon}}+v_{\varepsilon}\quad\mbox{ where }$ (4)
@@
$-\Delta u=u^{\frac{n+2}{n-2}},\,u>0\,\quad\mbox{ in }\quad\mathbb{R}^{n}.$
@@
Let $n\geq 4$ and $0<V\in C^{3}(\overline{\Omega})$ be a positive function and let $(u_{\varepsilon})$ be a sequence of solutions of $(\mathcal{P}_{\varepsilon,V})$ having the form (4) and satisfying (5), (6) and (8). Then, the following results hold:
@@
      or #$\#B(z)=1$, let $B(z)=\{a_{i,\varepsilon}\}$, it holds that $\beta(\varepsilon)|a_{i,\varepsilon}-z|$ is uniformly bounded where
@@
      or #$\#B(z)\geq 2$, let $B(z)=\{a_{i_{1},\varepsilon},\cdots,a_{i_{m},\varepsilon}\}$, it holds that $|\ln\varepsilon|^{\sigma_{n}/4}\varepsilon^{-(n-4)/(2n)}|a_{i_{j},\varepsilon}-z|$ is uniformly bounded for each $j\in\{1,\cdots,m\}$. Furthermore, there exist at least $m-1$ points $a_{i_{j},\varepsilon}$’s such that $|\ln\varepsilon|^{\sigma_{n}/4}\varepsilon^{-(n-4)/(2n)}|a_{i_{j},\varepsilon}-z|$ is uniformly bounded below. In this case $z$ is a non simple blow up point for the family $(u_{\varepsilon})$.
@@
The proof of Theorem 1.1 is based on some balancing conditions satisfied by the concentration parameters, that is, relationships that ensure the equilibrium between different factors governing the blow-up behavior of the solutions. These conditions are derived by performing an asymptotic expansion of the gradient of the functional $I_{\varepsilon,V}$ and testing the equation with suitable vector fields, which leads to constraints on the concentration points and the corresponding rates at which the solution blows up. Through a careful study of these conditions, we obtain a precise description of the blow-up phenomenon, including the locations and rates of concentration. Note that, in contrast to traditional asymptotic blow-up analysis techniques, which rely on precise pointwise $C^{0}$-estimates and extensive use of Pohozaev identities, our approach diverges from these methods. We believe that by eliminating the need for pointwise estimates and Pohozaev identities, our approach offers a powerful tool for addressing non-compact variational problems with more complex blow-up behavior. Indeed, the presence of non-simple blow-up points significantly complicates the establishment of pointwise $C^{0}$-estimates, posing a particularly challenging obstacle.
@@ 2 Balancing conditions of the concentration parameters
$u_{\varepsilon}=\sum_{i=1}^{N}\alpha_{i,\varepsilon}\pi\delta_{a_{i,\varepsilon},\lambda_{i,\varepsilon}}+v_{\varepsilon},\qquad\mbox{ with }$ (13)
@@
$E_{a,\lambda}:=\mbox{span}\Big{\{}\pi\delta_{a_{i},\lambda_{i}},\frac{\partial(\pi\delta_{a_{i},\lambda_{i}})}{\partial\lambda_{i}},\frac{\partial(\pi\delta_{a_{i},\lambda_{i}})}{\partial a_{i,j}},\,\,i\in\{1,\cdots,N\},\,\,j\in\{1,\cdots,n\}\Big{\}}.$ (15)
@@
$\alpha_{i}\|\pi\delta_{a_{i},\lambda_{i}}\|^{2}+\sum_{j\neq i}\alpha_{j}\langle\pi\delta_{a_{j},\lambda_{j}},\pi\delta_{a_{i},\lambda_{i}}\rangle=\int_{\Omega}u_{\varepsilon}^{p-\varepsilon}\pi\delta_{a_{i},\lambda_{i}}.$
@@
$\int_{\Omega}u_{\varepsilon}^{p-\varepsilon}\pi\delta_{a_{i},\lambda_{i}}=\alpha_{i}^{p-\varepsilon}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})^{p-\varepsilon+1}+O\Big{(}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})^{p-\varepsilon}\Big{[}\sum_{j\neq i}(\pi\delta_{a_{j},\lambda_{j}})+|v_{\varepsilon}|\Big{]}\Big{)}$
$+O\Big{(}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})\Big{[}\sum_{j\neq i}(\pi\delta_{a_{j},\lambda_{j}})^{p-\varepsilon}+|v_{\varepsilon}|^{p-\varepsilon}\Big{]}\Big{)}.$
@@
$(EAl_{i})\quad\left(1-\alpha_{i}^{p-1-\varepsilon}\lambda_{i}^{-\varepsilon(n-2)/2}\right)=O(R_{\alpha_{i}})\qquad\mbox{ with }R_{\alpha_{i}}:=\varepsilon+\frac{\ln^{\sigma_{n}}(\lambda_{i})}{\lambda_{i}^{2}}+\left\|v\right\|^{2}+\sum_{i\neq j}\varepsilon_{ij},$
@@
[…]

### v2

@@
Abstract: Consider a smooth, bounded domain $\Omega\subset\mathbb{R}^{n}$ with $n\geq 4$ and a smooth positive function $V$. We analyze the asymptotic behavior of a sequence of positive solutions $u_{\varepsilon}$ to the equation $\Delta u+V(x)u=u^{\frac{n+2}{n-2}-\varepsilon}$ in $\Omega$ with zero Dirichlet boundary conditions, as $\varepsilon\to 0$. We determine the precise blow-up rate and characterize the locations of interior concentration points in the general case of multiple blow-up, providing an exhaustive description of interior blow-up phenomena of this equation. Our result is established through a delicate analysis of the gradient of the corresponding Euler-Lagrange functional.
@@ 1 Introduction and Main Results
where $\Omega$ is a smooth bounded domain in $\mathbb{R}^{n}$, $n\geq 4$, the potential $V$ is a $C^{3}$ positive function on $\overline{\Omega}$, p+1=(2n)/(n2)p+1=(2n)/(n-2) is the critical Sobolev exponent for the embedding $H^{1}_{0}(\Omega)\hookrightarrow L^{q}(\Omega)$ and $\varepsilon$ is a small positive parameter.
@@
Over the past decades, a substantial body of literature has developed on the blow-up phenomena of solutions to semilinear equations with critical exponents, a topic too broad to summarize fully. Here, we highlight only a selection of some works, from which we hope a more comprehensive bibliography can be constructed. In the case where $V\equiv 0$, the asymptotic behavior of solutions that blow up at a single point has been well-established in [12, 38, 26]. The scenario of multiple concentration points, still for $V\equiv 0$, was later investigated in [39, 7, 40]. For $V\not\equiv 0$, single-blow-up behavior in the interior has been studied in the three-dimensional case by [25]. The focus of this paper is on multiple blow-up points in the interior for dimensions greater than or equal to 4. Other blow-up scenarios have also been studied, including those for sign-changing solutions and for positive solutions under Neumann boundary conditions. For additional references, we refer the reader to [16, 28, 29, 21, 24, 32, 1, 2, 30, 14, 27] and the works cited therein. For the inverse problem of constructing blow-up solutions in this setting, we direct the reader to [3, 10, 8, 7, 22, 11, 33, 41, 42, 18, 34, 36] and their respective references.
@@
defined on $H^{1}_{0}(\Omega)$ equipped with the norm $\|.\|$ and its corresponding inner product given by
@@
Since $V$ is a continuous positive function on $\overline{\Omega}$, we see that this norm is equivalent to the two norms $||.||_{0}$ and $||.||_{1}$ of $H^{1}_{0}(\Omega)$ and $H^{1}(\Omega)$ respectively.
Note that all solutions $u_{\varepsilon}$ of $(\mathcal{P}_{\varepsilon,V})$ satisfy $\|u_{\varepsilon}\|\geq C$ with a positive constant $C$ independent of the parameter $\varepsilon$. Therefore, the concentration compactness principle ([43, 31]) implies that if $u_{\varepsilon}$ is an energy-bounded solution of $(\mathcal{P}_{\varepsilon,V})$ that weakly ( but not strongly) converges to $\omega$ (where $\omega=0$ or $\omega$ is a solution of $(\mathcal{P}_{0,V})$), then $u_{\varepsilon}$ must concentrate at a finite number $N$ of points in $\overline{\Omega}$. More precisely, $u_{\varepsilon}$ can be expressed as:
$u_{\varepsilon}=\omega+\sum_{i=1}^{N}\delta_{{a_{i,\varepsilon}},\lambda_{i,\varepsilon}}+v_{\varepsilon}\quad\mbox{ where }$ (4)
@@
$\Delta u=u^{\frac{n+2}{n-2}},\,u>0\,\quad\mbox{ in }\quad\mathbb{R}^{n}.$
It is worth noting that the following related problem was investigated in [28] for the case $n\geq 4$
$(\mathcal{Q}_{\varepsilon,V}):\qquad\begin{cases}-\Delta u+\varepsilon Vu=u^{p},\quad u>0\quad&\mbox{in}\quad\Omega,\\ \quad u=0\,\quad&\mbox{on}\quad\partial\Omega.\end{cases}$
In [28], by considering a blowing-up family of solutions $(u_{\varepsilon})$, the authors proved that when $V<0$, the concentration points $a_{i,\varepsilon}$’s remain away from both the boundary and each other, that is, $d(a_{i,\varepsilon},\partial\Omega)\geq c>0$ and $|a_{i,\varepsilon}-a_{j,\varepsilon}|\geq c>0$ for each $i\neq j$. In contrast, for our problem, $(\mathcal{P}_{\varepsilon,V})$, the behavior is fundamentally different. We believe that when $V<0$ and $\Delta+V$ is coercive, blow-up phenomena do not occur, whereas they may arise when $V>0$. Moreover, some concentration points may converge to the same location (see the theorems below). Additionally, it appears possible to construct solutions featuring concentration points -either isolated or clustered- located on the boundary, see the end of Subsection 2.1. We plan to explore this question in future work.
@@
Let $n\geq 4$ and $0<V\in C^{3}(\overline{\Omega})$ be a positive function and let $(u_{\varepsilon})$ be a sequence of solutions of $(\mathcal{P}_{\varepsilon,V})$ having the form (4), with $\omega=0$ (i.e. $u_{\varepsilon}$ converges weakly to 0), and satisfying (5), (6) and (8). Then, the following results hold:
@@
      or #B(z)=1\#B(z)=1, let $B(z)=\{a_{i,\varepsilon}\}$, it holds that $\beta(\varepsilon)|a_{i,\varepsilon}-z|$ is uniformly bounded where
@@
      or $B(z)\geq 2$, let $B(z)=\{a_{i_{1},\varepsilon},\cdots,a_{i_{m},\varepsilon}\}$, it holds that $|\ln\varepsilon|^{\sigma_{n}/4}\varepsilon^{-(n-4)/(2n)}|a_{i_{j},\varepsilon}-z|$ is uniformly bounded for each $j\in\{1,\cdots,m\}$. Furthermore, there exist at least m1m-1 points $a_{i_{j},\varepsilon}$’s such that $|\ln\varepsilon|^{\sigma_{n}/4}\varepsilon^{-(n-4)/(2n)}|a_{i_{j},\varepsilon}-z|$ is uniformly bounded below. In this case $z$ is a non simple blow up point for the family $(u_{\varepsilon})$.
@@
The next theorem provides a description of the interior blow-up phenomenon occurring when the sequence of solutions converges weakly to a non-zero limit.
Theorem 1.2
Let $n\geq 7$ and $0<V\in C^{3}(\overline{\Omega})$ be a positive function and let $(u_{\varepsilon})$ be a sequence of solutions of $(\mathcal{P}_{\varepsilon,V})$ having the form (4) with $\omega\neq 0$ and satisfying (5), (6) and (8). Then, the conclusions of Theorem 1.1 hold with, in case $(ii)$, $\beta(\varepsilon)=-1/2$ for $n\geq 8$ and $\beta(\varepsilon)=-1/4$ for $n=7$.
Remark 1.3
In lower dimensions, that is, for $3\leq n\leq 5$, we believe that $(\mathcal{P}_{\varepsilon,V})$ admits no interior blowing up solutions whose weak limit is a non-zero solution of the limit problem $(\mathcal{P}_{0,V})$, see $(2)$ in Remark 2.6. This question will be addressed in a future work.
The proofs of our results are based on some balancing conditions satisfied by the concentration parameters, that is, relationships that ensure the equilibrium between different factors governing the blow-up behavior of the solutions. These conditions are derived by performing an asymptotic expansion of the gradient of the functional $I_{\varepsilon,V}$ and testing the equation with suitable vector fields, which leads to constraints on the concentration points and the corresponding rates at which the solution blows up. Through a careful study of these conditions, we obtain a precise description of the blow-up phenomenon, including the locations and rates of concentration.
@@ 2 Balancing conditions of the concentration parameters
$u_{\varepsilon}=\alpha_{0,\varepsilon}\omega+\sum_{i=1}^{N}\alpha_{i,\varepsilon}\pi\delta_{a_{i,\varepsilon},\lambda_{i,\varepsilon}}+v_{\varepsilon},\qquad\mbox{ with }$ (13)
@@
$E_{a,\lambda}:=\mbox{span}\Big{\{}\omega,\pi\delta_{a_{i},\lambda_{i}},\frac{\partial(\pi\delta_{a_{i},\lambda_{i}})}{\partial\lambda_{i}},\frac{\partial(\pi\delta_{a_{i},\lambda_{i}})}{\partial a_{i,j}},\,\,i\in\{1,\cdots,N\},\,\,j\in\{1,\cdots,n\}\Big{\}}.$ (15)
@@
$\alpha_{i}\|\pi\delta_{a_{i},\lambda_{i}}\|^{2}+\sum_{j\neq i}\alpha_{j}\langle\pi\delta_{a_{j},\lambda_{j}},\pi\delta_{a_{i},\lambda_{i}}\rangle+\alpha_{0}\langle\omega,\pi\delta_{a_{i},\lambda_{i}}\rangle=\int_{\Omega}u_{\varepsilon}^{p-\varepsilon}\pi\delta_{a_{i},\lambda_{i}}.$
Observe that
$\langle\omega,\pi\delta_{a_{i},\lambda_{i}}\rangle=\int_{\Omega}\delta_{a_{i},\lambda_{i}}^{(n+2)/(n-2)}\omega=\overline{c}_{1}\frac{\omega(a_{i})}{\lambda_{i}^{(n-2)/2}}+O\Big{(}\frac{\ln\lambda_{i}}{\lambda_{i}^{(n+2)/2}}\Big{)}$ (16)
with
$\overline{c}_{1}:=c_{0}^{\frac{n+2}{n-2}}\int_{\mathbb{R}^{n}}\frac{1}{(1+|x|^{2})^{(n+2)/2}}dx.$
@@
$\int_{\Omega}u_{\varepsilon}^{p-\varepsilon}\pi\delta_{a_{i},\lambda_{i}}=\alpha_{i}^{p-\varepsilon}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})^{p-\varepsilon+1}+O\Big{(}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})^{p-\varepsilon}\Big{[}\omega+\sum_{j\neq i}(\pi\delta_{a_{j},\lambda_{j}})+|v_{\varepsilon}|\Big{]}\Big{)}$
$+O\Big{(}\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})\Big{[}\omega^{p}+\sum_{j\neq i}(\pi\delta_{a_{j},\lambda_{j}})^{p-\varepsilon}+|v_{\varepsilon}|^{p-\varepsilon}\Big{]}\Big{)}.$
@@
$\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})^{p-\varepsilon}\omega+\int_{\Omega}(\pi\delta_{a_{i},\lambda_{i}})\omega^{p}\leq c\int_{\Omega}[\delta_{a_{i},\lambda_{i}}+\delta_{a_{i},\lambda_{i}}^{p}]=O\Big{(}\frac{1}{\lambda_{i}^{(n-2)/2}}\Big{)}=o(1),$ (19)
@@
2.1 Case of zero weak limit
In this subsection, we will assume that $\omega=0$.
@@ 2.1 Case of zero weak limit
$(EAl_{i})\quad\left(1-\alpha_{i}^{p-1-\varepsilon}\lambda_{i}^{-\varepsilon(n-2)/2}\right)=O(R_{\alpha_{i}})\qquad\mbox{ with }R_{\alpha_{i}}:=\varepsilon+\frac{\ln^{\sigma_{n}}(\lambda_{i})}{\lambda_{i}^{2}}+\left\|v\right\|^{2}+\sum_{i\neq j}\varepsilon_{ij},$ (22)
@@
$c_{2}:=\Big{(}\frac{n-2}{2}\Big{)}^{2}c_{0}^{\frac{2n}{n-2}}\int_{\mathbb{R}^{n}}\frac{(\left|x\right|^{2}-1)}{(1+\left|x\right|^{2})^{n+1}}\ln(1+\left|x\right|^{2})dx>0\,,\quad\overline{c}_{2}:=c_{0}^{\frac{2n}{n-2}}\int_{\mathbb{R}^{n}}\frac{1}{(1+|x|^{2})^{(n+2)/2}}dx,$